OpenStudy (anonymous):

a basket contains 6 pieaces of paper numbered from 1 to 6. a hat contains the letter a,b,and c. if a blindfolded person picks a number and a letter,what is the probability that the number is even and the letter is a

OpenStudy (anonymous):

Well what is the probability that the number is even?

OpenStudy (amistre64):

the probability of event1 AND event2 = p(event1) * p(event2)

OpenStudy (amistre64):

since the basket and hat are disjoint; theres no need to confuse yourself over the probabilty event2 given event1 :)

OpenStudy (anonymous):

6 pieaces =1,2,3,4,5,6, even = 2,4,6 3/6=1/2

OpenStudy (amistre64):

yes; 1/2 for the first event; and for the next?

OpenStudy (anonymous):

I dont know next, becaause letter I don't get what is say

OpenStudy (anonymous):

amistree u do next I don't know next

OpenStudy (amistre64):

there are 3 total choices and 1 is an a :) 1/3

OpenStudy (amistre64):

P(even#) * P(a) = (1/2) * (1/3) = (1/6)